a person A of 50 kg rest on a swing length 1m making an angle 37° with vertical another person But pushes him to swing on the other side 53° with vertical . the work done by the person B is

Grade:11

2 Answers

Ankita
5 years ago
100 joules must be the answre.. just apply work energy theorem properly to get it..........................😎😎😎
Ankita
5 years ago
Draw the scene on paper to figure out that there is a net displacement of 1/5 m in y axis .
Sowork done by gravity is m×g× 1/5= 💯 joules😎😎...
 
Displacement =1/5 because 2/5. -1/5=1/5 ..
And i don't know how to add attachment 🤣🤣🤣🤣😎😎😎 
 
100joules😎😎
  1.